OpenStudy (anonymous):

Quick question: Why can't you factor the "sum of two squares"? Please make it simple... ^-^ Thanks!! (I'm just curious.)

OpenStudy (anonymous):

Because you have to have that third term of: 2xy That's if you wanted to factor: x^2 + y^2 You can't because you need that 2xy : x^2 + 2xy + y^2 = (x + y)(x + y) = (x + y)^2

OpenStudy (anonymous):

All good now, @ashizepic ?

OpenStudy (anonymous):

You would otherwise have to start with: (x + y)(x + y) or (x - y)(x - y) to keep that sign positive in "the sum of 2 squares" and either of these produces a third term. Unlike: (x + y)(x - y) = x^2 - y^2 where you can have the difference of 2 squares because here, that third term vanishes. That's a little bit more explanation. But you wanted it short initially.

OpenStudy (amistre64):

you can factor the sum of 2 squares ... but it depends on what type of solutions you are looking for

OpenStudy (anonymous):

Yes, if you allow complex numbers.

OpenStudy (anonymous):

Ohhh... thank you!

OpenStudy (anonymous):

uw!